c. What will be the new bottleneck in the process?

d . How many customers a day will not be served? Which customers are going to wait? Explain your reasoning!   

(Car Wash Supply Process) CC Car Wash specializes in car cleaning services. The ser- vices offered by the company, the exact service time, and the resources needed for each of them are described in the table following:
Resource Used Service Description Processing Time A. Wash Exterior car washing and drying 10 min 1 automated washing machine 10 min B. Wax Exterior car waxing 1 automated waxing machine 7 min C. Wheel cleaning Detailed cleaning of all wheels 1 employee D. Interior cleaning Detailed cleaning inside the car 1 employee 20 min The company offers the following packages to their customers: Package 1: Includes only car wash (service A) Package 2: Includes car wash and waxing (services A and B) Package 3: Car wash, waxing, and wheel cleaning (services A, B, and C) . Package 4: All four services (A, B, C, and D) Customers of CC Car Wash visit the station at a constant rate (you can ignore any effects of variability) of 40 customers per day. Of these customers, 40 percent buy Package 1, 15 percent buy Package 2, 15 percent buy Package 3, and 30 percent buy Package 4. The mix does not change over the course of the day. The store operates 12 hours a day. a. What is the implied utilization of the employee doing the wheel cleaning service? b. Which resource has the highest implied utilization? For the next summer, CC Car Wash anticipates an increase in the demand to 80 cus- tomers per day. Together with this demand increase, there is expected to be a change in the mix of packages demanded: 30 percent of the customers ask for Package 1, 10 percent for Package 2, 10 percent for Package 3, and 50 percent for Package 4. The company will install an additional washing machine to do service A c. What will be the new bottleneck in the process? d. How many customers a day will not be served? Which customers are going to wait? Explain your reasoning!
Answer #1

(c)

Resource Demand (inc. per day) Capacity (inc. per day) Utilization
Wash 80 2*(60/12)*12 = 120 80/120 = 67%
Wax 80*(0.10+0.10+0.50) = 56 (60/12)*12 = 60 56/60 = 93%
Wheel 80*(0.10+0.50) = 48 (60/8)*12 = 90 48/90 = 53%
Interior 80*(0.50) = 40 (60/15)*12 = 48 40/48 = 83%

The wax resource or the wax machine as per the above calculations has the highest percentage of utilization, so wax machine will be considered as bottleneck in this.